OpenStudy (mathmath333):

find the average increase rate if the increase in population for first year is 30 percent and that in the second year is 40 percent

OpenStudy (anonymous):

The average would be 35%

OpenStudy (mathmath333):

a.41 b. 40 c. 56 d. 38

OpenStudy (mathmath333):

@surjithayer

OpenStudy (amistre64):

assume we start with 1 1(1.3) is the value of the first time 1(1.3)(1.4) is the value after the second time .... right??

OpenStudy (mathmath333):

ok

OpenStudy (amistre64):

so, what percentage does it take for 2 periods to go from 1 to 1(1.3)(1.4)?

OpenStudy (amistre64):

1(1.r)(1.r) = 1(1.3)(1.4)

OpenStudy (mathmath333):

70 ?

OpenStudy (amistre64):

(1+r)^2 = 1.3(1.4) = 1.82 so in 2 years we have a .82 increase

OpenStudy (amistre64):

the 1+r may be extra baggage

OpenStudy (mathmath333):

1(1.3)(1.4) why r u multiplying 1 , 1.3 and 1.4

OpenStudy (amistre64):

spose we start with a population of 1 after the first time period we have 1(1.3) since we are told it has a 30$ increase our new population is 1(1.3) and we increase by 40%: 1(1.3)(1.4) = 1.82 overall right?

OpenStudy (amistre64):

other than that, what doe you material suggest you do to determine a solution? i may just be not reading this right

OpenStudy (mathmath333):

its from the chapter Averages

OpenStudy (mathmath333):

it gives answer 41

OpenStudy (amistre64):

oh good, then my idea has some merit :)

OpenStudy (amistre64):

.82 in 2 years is an average of .82/2

OpenStudy (amistre64):

if we assume for simplicity that the original population is 1 year0= 1 year1=1(1.3) = 1.3 year2=1.3(1.4) = 1.82 1.82 - 1= .82 increase over 2 years

OpenStudy (mathmath333):

book gave this one ->100 130 182 i got it 30 52 =(30+52)/2=41

OpenStudy (amistre64):

same idea yes